OpenStudy (anonymous):

Can someone please help for a fan and medal?!?!?! Evaluate. cube root of 7 multiplied by the square root of 7 over sixth root of 7 to the power of 5

OpenStudy (anonymous):

\[7^{\frac{ 1 }{ 3 }} \times 7^{\frac{ 1 }{ 2 }} \times ?\] i didnt get qustion correctly

OpenStudy (anonymous):

This is what it looks like:

OpenStudy (anonymous):

Convert all of the roots into exponents. Then, you can use exponent rules to simplify it.

OpenStudy (anonymous):

I'm sorry, but can you tell me what the roots are? It's been a while since I did this so I don't know what parts of the problem are called.

OpenStudy (anonymous):

All I remember is what the radical is.

OpenStudy (anonymous):

\[\sqrt[2]{7}=7^\frac{ 1 }{ 2 }\] \[\sqrt[3]{7}=7^\frac{ 1 }{ 3 }\] and so on

OpenStudy (anonymous):

If the number inside the root is to a degree, then it would be like this: \[\sqrt[a]{x^b}=x^\frac{ b }{ a}\]

OpenStudy (anonymous):

Okay, thank you! I can continue from here.

OpenStudy (anonymous):

Hey @greenglasses would the answer be \[7\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can you show me what you did? I'm getting 1.

OpenStudy (anonymous):

Hang on, let me look over what I did.

OpenStudy (anonymous):

Oooooohh I see what I did wrong. Okay, I did the work again and I got 1 also. No need for you to waste your time going over my work. I made a bad mistake that I fixed.

OpenStudy (anonymous):

Alright, if you've got it that's good then :)

OpenStudy (anonymous):

Thank you again!